# Topological characterization of phase transitions and critical edge states in one-dimensional non-Hermitian systems with sublattice symmetry

### v1

@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grants No. 12275260 and No. 11905211), the Fundamental Research Funds for the Central Universities (Grant No. 202364008), and the Young Talents Project of Ocean University of China.
Appendix A Calculation of the edge states
@@
In lattice representation, the Hamiltonian of NHSSH $(0,1)$ chain in Sec. III.3 takes the form
@@
Following Ref. [25], we apply the similarity transformation $\hat{S}$ to $\hat{H}_{01}$ under the OBC, where
@@
We choose $\varrho$ to ensure the left and right intracell hoppings to have equal magnitudes, which means that $|\rho|=\sqrt{|J+\gamma|/|J-\gamma|}$. Let us consider a chain with unit cell indices n=1,2,,Nn=1,2,...,N and take the OBC at both ends. A zero-energy eigenmode of $\hat{H}^{\prime}_{01}$ satisfies $\hat{H}^{\prime}_{01}|\psi_{0}\rangle=0$. Expanding $|\psi_{0}\rangle$ in the lattice representation as $|\psi_{0}\rangle=\sum_{m}\left(A_{m}\hat{a}_{m}^{\dagger}+B_{m}\hat{b}_{m}^{\dagger}\right)|\emptyset\rangle$ ($|\emptyset\rangle$ denotes the vacuum state) and applying $\hat{H}^{\prime}_{01}$ to it, we arrive at the iteration equations of wave amplitudes $(A_{n},B_{n})$, i.e.,
$\varrho^{-1}(J+\gamma)B_{1}=0,\quad\varrho(J-\gamma)A_{N}=0,$ (39)
$\varrho(J-\gamma)A_{n}+A_{n+1}=0,\quad n=1,...,N-1,$ (40)
$\varrho^{-1}(J+\gamma)B_{n+1}+B_{n}=0,\quad n=1,...,N-1.$ (41)
@@
$|\psi_{0}^{{\rm R}}\rangle=$ $\sum_{n=1}^{N}[-\varrho^{-1}(J+\gamma)]^{N-n}\hat{b}_{n}^{\dagger}|\emptyset\rangle.$ (42)
It is clear that the states $|\psi_{0}^{{\rm L}}\rangle$ and $|\psi_{0}^{{\rm R}}\rangle$ represent a left-localized and right-localized edge zero modes if and only if $|\varrho(J-\gamma)|<1$ and $|(J+\gamma)/\varrho|<1$, respectively. Both these conditions yield $|J^{2}-\gamma^{2}|<1$, which is exactly the condition for us to have topologically nontrivial gapped phases with the winding number $W=1$ and the number of zero-energy edge modes $N_{0}=2$, as reported in Eqs. (25) and (26). The bulk-edge correspondence of our NHSSH $(0,1)$ chain is thus proved. Importantly, we find no zero-energy edge modes when $J^{2}-\gamma^{2}=\pm 1$, i.e., at the critical points of the system under the OBC. This observation confirms that the gapless phase boundary of the NHSSH $(0,1)$ chain is indeed topologically trivial from the perspective of critical edge states.
We next consider the NHSSH $(1,2)$ chain in Sec. III.3, whose Hamiltonian takes the form
@@
Under the similarity transformation
@@
$|\varphi_{0}^{{\rm L,1}}\rangle=\hat{a}_{1}^{\dagger}|\emptyset\rangle,\quad|\varphi_{0}^{{\rm R,1}}\rangle=\hat{b}_{N}^{\dagger}|\emptyset\rangle,$ (46)
which form a pair of degenerate edge modes at the left and right ends of the lattice. To find other possible solutions, we again expand the zero mode in the lattice space as $|\varphi_{0}\rangle=\sum_{m}\left(C_{m}\hat{a}_{m}^{\dagger}+D_{m}\hat{b}_{m}^{\dagger}\right)|\emptyset\rangle$ and apply the Hamiltonian $\hat{H}^{\prime}_{12}$ to it. The resulting zero-energy solutions in the limit $N\rightarrow\infty$ are given by
@@
$|\varphi_{0}^{{\rm R,2}}\rangle=$ $\sum_{n=1}^{N-1}[-\varrho^{-1}(J+\gamma)]^{N-n-1}\hat{b}_{n}^{\dagger}|\emptyset\rangle.$ (47)

### v2

@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grants No. 12275260 and No. 11905211), the Fundamental Research Funds for the Central Universities (Grant No. 202364008), and the Young Talents Project of Ocean University of China. Note added.—After the submission of this manuscript, we noted a preprint arXiv:2509.09587 on non-Hermitian criticality enriched by PT-symmetry, where the bulk theory is deduced with respect to the conventional Brillouin zone of Bloch bands. In our case, the non-Hermitian topological criticality is protected by sublattice symmetry, and the theory is developed with respect to the generalized Brillouin zone of non-Bloch bands, which is also reducible to Bloch band cases in Hermitian limits.
Appendix A Theory: further details
In this Appendix, we offer additional details and illustrative examples about our theory as sketched in Sec. II.
We first provide some further analyses of the non-Hermitian Hamiltonian in Eq. (1). The off-diagonal elements $f(k)$ and $g(k)$ in Eq. (1) are both $2\pi$-periodic in $k$, and the $H(k)$ is non-Hermitian if there exists a quasimomentum $k$ such that $f(k)\neq g^{*}(k)$. The sublattice symmetry of $H(k)$ is given by $\Gamma=\sigma_{z}$, in the sense that $\Gamma H(k)\Gamma=-H(k)$ and $\Gamma^{2}=\sigma_{0}$, where $\sigma_{0}$ denotes the $2\times 2$ identity matrix. The Hamiltonian $H(k)$ has two energy bands, whose dispersion relations are given by
$E_{\pm}(k)\equiv\pm E(k)=\pm\sqrt{f(k)g(k)}.$ (29)
$\frac{b(k)}{a(k)}=\frac{E(k)}{f(k)}=\frac{g(k)}{E(k)}.$ (30)
$P(k)\equiv{g(k)}/{f(k)},\qquad Q(k)\equiv f(k)g(k).$ (31)
As the functions $f(k)$ and $g(k)$ are both $2\pi$-periodic in $k$, the $P(k)$ and $Q(k)$ in Eq. (31) should also satisfy $P(k+2\pi)=P(k)$ and $Q(k+2\pi)=Q(k)$. We could thus expand the $P(k)$ and $Q(k)$ into Fourier series as $P(k)=\sum_{n\in\mathbb{Z}}p_{n}e^{ink}$ and $Q(k)=\sum_{n\in\mathbb{Z}}q_{n}e^{ink}$. These series can be viewed as polynomials of the exponential factor $e^{ik}$. To treat Bloch and non-Bloch band theories of non-Hermitian systems on an equal footing, we extend the $e^{ik}$ to the whole complex plane by setting $e^{ik}\rightarrow z$ with $z\in\mathbb{C}$. The resulting characteristic polynomials of the complex-extended Hamiltonian $H(z)$ are given by $P(z)=\sum_{n\in\mathbb{Z}}p_{n}z^{n}$ and $Q(z)=\sum_{n\in\mathbb{Z}}q_{n}z^{n}$, or Eq. (2) in the main text.
In Sec. II, we have mentioned the possible issue that previous theories may encounter at gapless critical points. The issue can be made clear via a counterexample. Let us consider a complex-continued non-Hermitian Hamiltonian $H(z)$ with off-diagonal elements $f(z)=J_{0}+J_{1}z^{-1}$ and $g(z)=J_{0}+J_{1}z$. When $J_{0}$ and $J_{1}$ take real values, we go back to the standard SSH model after setting $z\rightarrow e^{ik}$. Following the $P(z)$-based approach as discussed in Sec. II, we have $m=1$, $n=0$ and thus $2\mu=2\max\{1,0\}=2$. By definition, we find $P(z)=g(z)/f(z)=z(J_{0}+J_{1}z)/(J_{0}z+J_{1})$. It can then be identified that if $|J_{1}|<|J_{0}|$, the first $2$ zeros/poles of $P(z)$ contains a zero at $z=0$ and a pole at $z=-J_{1}/J_{0}$, yielding $N_{z}^{\in 2}=1$, $N_{p}^{\in 2}=1$, $W_{0}=N_{z}^{\in 2\mu}-N_{p}^{\in 2\mu}=0$ and $N_{0}=|W_{0}|=0$. Meanwhile, if $|J_{1}|>|J_{0}|$, the first $2$ zeros/poles of $P(z)$ contains two zeros at $z=0$ and $z=-J_{0}/J_{1}$, yielding $N_{z}^{\in 2}=2$, $N_{p}^{\in 2}=0$, $W_{0}=N_{z}^{\in 2\mu}-N_{p}^{\in 2\mu}=2$ and $N_{0}=|W_{0}|=2$. These predictions are all correct, as the two bulk bands of $H(z)$ are gapped at zero energy $Q(z)\neq 0$] for any finite $z$ whenever $|J_{1}|\neq|J_{0}|$. However, at the phase transition point $|J_{1}|=|J_{0}|$, we have $|-J_{1}/J_{0}|=|-J_{0}/J_{1}|=1$, which means that we could not order the magnitudes of the zero and pole at $z=-J_{0}/J_{1}$ and $z=-J_{1}/J_{0}$. The determination of which are the first two zeros and/or poles of $P(z)$ then becomes ambiguous. If we include or exclude both their contributions to $W_{0}$, we will get $W_{0}=1$ and thus $N_{0}=1$. This is impossible, as the zero modes must come in pairs (i.e., $N_{0}\in 2\mathbb{N}$) due to the sublattice symmetry. If we only retain the zero at $z=-J_{0}/J_{1}$ as the second zero/pole, we will get $W_{0}=2$ and $N_{0}=2$, which is incorrect even for the Hermitian SSH model [46]. If we instead retain the pole at $z=-J_{1}/J_{0}$ as the second zero/pole, we will find $W_{0}=0$ and $N_{0}=0$. Even though these are the correct values of topological invariants and edge states at the critical point of SSH model [46], they could not be conclusively and unambiguously predicted by the $P(z)$-based approach. Therefore, the algebraic scheme in Refs. [27, 30] tends out to be inapplicable to the characterization of nontrivial topology and bulk-edge correspondence at the phase transition (bulk-gap closing) point of non-Hermitian systems. Similarly, as the algebraic and GBZ schemes are proved to be equivalent for sublattice-symmetric models [30], the standard GBZ approach is also expected to be unworkable for describing gSPT phases in non-Hermitian systems.
These observations motivate us to generalize existing theories in order to characterize the nontrivial topology at non-Hermitian critical points. The above analysis suggests that for the SSH model, there could be competing zeros at $z=-J_{0}/J_{1}$ and poles at $z=-J_{1}/J_{0}$ of equal magnitudes in $P(z)$ at the critical point. Their contributions are encoded in $Q(z)$ as its zeros at the gapless point. Therefore, we introduce another winding number by applying the Cauchy’s argument principle to $Q(z)$. The winding numbers of $P(z)$ and $Q(z)$ are finally given by Eq. (3) of the main text.
We may now check the applicability of our theory to the SSH model. Besides the $P(z)=z(J_{0}+J_{1}z)/(J_{0}z+J_{1})$, we also have $Q(z)=(J_{0}z+J_{1})(J_{0}+J_{1}z)/z$. The GBZ here is just the standard Brillouin zone, which forms a circle of radius $1$ on the complex plane regardless of whether the values of $J_{0}$ and $J_{1}$ are real or complex. Direct calculations following Eqs. (3)–(5) in the main text predict $(W_{1},W_{2})=(0,0)$ when $|J_{1}|<|J_{0}|$, $(W_{1},W_{2})=(2,0)$ when $|J_{1}|>|J_{0}|$, and $(W_{1},W_{2})=(1,-1)$ when $|J_{1}|=|J_{0}|$, yielding $W=0,1,0$ and $N_{0}=0,2,0$ in these three parameter regions. Besides reproducing known results about gapped phases [97], we are now able to confirm the topological triviality of the gapless phase boundary $|J_{1}|=|J_{0}|$ of the SSH model [47] without ambiguity.
Appendix B Single $\alpha$ chain: gapped and gapless phases
In this Appendix, we investigate the simplest type of NHSSH $\alpha$ chain, whose Hamiltonian is given by ${\hat{H}}={\hat{H}}_{\alpha}$ for a fixed $\alpha$. This NHSSH single $\alpha$ chain has the complex-extended Hamiltonian $H(z)=H_{\alpha}(z)$. It contains only a single length scale in its hopping amplitudes $J_{\alpha}^{{\rm L,R}}$. Throughout this Appendix, we assume $\alpha\geq 0$ without losing generality.
$P(z)=\frac{J_{\alpha}^{{\rm L}}}{J_{\alpha}^{{\rm R}}}z^{2\alpha},\qquad Q(z)=J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}.$ (32)
Since the $Q(z)=[E(z)]^{2}$ is independent of $z$, the bulk spectrum of the system is formed by two flat bands at $\pm E(z)=\pm\sqrt{J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}}$, and the GBZ is just the standard BZ with $k\in[-\pi,\pi]$. On the complex $z$-plane, the $Q(z)$ has neither zeros nor poles, while the $P(z)$ has a zero of order $2\alpha$ at $z=0$, which is inside the BZ unit circle $|z|=1$. According to the definition of winding numbers for $P(z)$ and $Q(z)$ in Eq. (3), we find $W_{1}=2\alpha$ and $W_{2}=0$. The topological invariant $W$ of the system and the number of its degenerate zero modes $N_{0}$ under the OBC are thus given by Eqs. (4) and (5) as
$W=\frac{W_{1}+W_{2}}{2}=\alpha,\quad N_{0}=2|W|=2\alpha.$ (33)
We now verify our results about the NHSSH single $\alpha$ chain with two groups of numerical examples. In the first group, we consider the case of asymmetric hopping with $(J_{\alpha}^{{\rm L}},J_{\alpha}^{{\rm R}})=(0.5,2)$ for $\alpha=0,1,2$. The bulk spectrum is expected to show two flat bands at $\pm\sqrt{J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}}=\pm 1$, separated by a constant gap $\Delta E=2$. In the second group, we consider the case of unidirectional hopping with $(J_{\alpha}^{{\rm L}},J_{\alpha}^{{\rm R}})=(0,1)$ for $\alpha=0,1,2$. The bulk spectrum is expected to show two degenerate flat bands at zero energy. In these cases, the theoretically predicted winding numbers and numbers of zero-energy edge modes are $W=(0,1,2)$ and $N_{0}=(0,2,4)$ according to Eq. (33), regardless of the chosen parameter values $(J_{\alpha}^{{\rm L}},J_{\alpha}^{{\rm R}})$.
Figure 10: Energy spectra $E$ and edge states of NHSSH single $\alpha$ chains with bulk gaps. The system has $10$ unit cells and 2020 lattice sites. $j$ is the state index. $l$ is the lattice index. (a) Spectrum of $\hat{H}_{0}$ with $(J_{0}^{{\rm L}},J_{0}^{{\rm R}})=(0.5,2)$ under the PBC (in circles) and OBC (in stars). (b) Spectrum of $\hat{H}_{1}$ with $(J_{1}^{{\rm L}},J_{1}^{{\rm R}})=(0.5,2)$ under the PBC (in circles) and OBC (in stars). (c) Spectrum of $\hat{H}_{2}$ with $(J_{2}^{{\rm L}},J_{2}^{{\rm R}})=(0.5,2)$ under the PBC (in circles) and OBC (in stars). (d) Probability distributions $P$ of the zero-energy edge states of $\hat{H}_{1}$ in (b). (e) Probability distributions $P$ of the zero-energy edge states of $\hat{H}_{2}$ in (c). In (d) and (e), different edge states are distinguished by the colors of probability bars.
Figure 11: Energy spectra $E$ and edge states of NHSSH single $\alpha$ chains without bulk gaps. The system has $10$ unit cells and 2020 lattice sites. $j$ is the state index. $l$ is the lattice index. (a) Spectrum of $\hat{H}_{0}$ with $(J_{0}^{{\rm L}},J_{0}^{{\rm R}})=(0,1)$ under the PBC (in circles) and OBC (in stars). (b) Spectrum of $\hat{H}_{1}$ with $(J_{1}^{{\rm L}},J_{1}^{{\rm R}})=(0,1)$ under the PBC (in circles) and OBC (in stars). (c) Spectrum of $\hat{H}_{2}$ with $(J_{2}^{{\rm L}},J_{2}^{{\rm R}})=(0,1)$ under the PBC (in circles) and OBC (in stars). (d) Probability distributions $P$ of the zero-energy edge states of $\hat{H}_{1}$ in (b). (e) Probability distributions $P$ of the zero-energy edge states of $\hat{H}_{2}$ in (c). In (d) and (e), different edge states are distinguished by the colors of probability bars.
In Fig. 11, we present the energy spectra and edge states of the single $\alpha$ chain for $\alpha=0,1,2$ and $(J_{\alpha}^{{\rm L}},J_{\alpha}^{{\rm R}})=(0,1)$. The spectrum of $\hat{H}_{0}$ in Fig. 11(a) has no edge zero modes ($N_{0}=0$). The spectrum of $\hat{H}_{1}$ in Fig. 11(b) holds two zero-energy edge modes ($N_{0}=2$), whose distributions are shown in Fig. 11(d). The spectrum of $\hat{H}_{2}$ in Fig. 11(c) contains four zero-energy edge modes ($N_{0}=4$), whose distributions are shown in Fig. 11(e). These results again verify our theoretical predictions. Notably, even though the bulk spectra of $\hat{H}_{0,1,2}$ have no gaps at $E=0$ in Figs. 11(a)–11(c), our theory could still correctly predict the numbers of edge zero modes that can coexist with gapless bulks in these cases. Note that for our models at half-filling, gaplessness means that the two bulk bands are touched at least at one point in the energy spectrum. This is regardless of whether the two bands are presented in momentum or lattice representation. Since the spectra in Figs. 11(a–c) are obtained under both the periodic and open boundary conditions in lattice representation, we do not have a conserved quasimomentum $k$ in both cases. The spectra are thus shown with respect to the index of energy eigenstate there.
The Figs. 11(a–c) are obtained for a special choice of system parameters, under which the two bulk bands of the system are completely degenerate at zero energy. One may also see this from the Bloch Hamiltonian of the bulk under periodic boundary condition, which in this case is given by $H_{\alpha}(k)=J_{\alpha}^{\rm R}(\sigma_{x}+i\sigma_{y})e^{-ik\alpha}/2$. It has the dispersion relation $E_{\pm}(k)=0$ for any $\alpha$. Therefore, the two bulk bands are both flat and they are degenerate at every $k\in[-\pi,\pi)$. The bulk spectrum of the system is further gapless in this case for any $J_{\alpha}^{\rm R}$, thus realizing a gapless phase in the parameter space $J_{\alpha}^{\rm R}\in{\mathbb{C}}$.
The presence or absence of edge zero modes in the spectrum of $\hat{H}_{\alpha}$ for a given $\alpha$ can be understood from the geometric connectivity of the underlying lattice [see Fig. 1(c)]. For $\hat{H}_{0}$, the sublattices A and B in each unit cell are coupled due to the non-vanishing $J_{0}^{{\rm L}}$ and/or $J_{0}^{{\rm R}}$, making it impossible to have isolated sites at boundaries to host edge zero modes. For $\hat{H}_{1}$, the sublattice B in unit cell $n$ is coupled to the sublattice A in unit cell $n+1$ for n=1,,N1n=1,...,N-1. The sublattices A in unit cell $1$ and B in unit cell $N$ are thus isolated from other sites, making them available to host two edge zero modes $\hat{a}_{1}^{\dagger}|\emptyset\rangle$ and $\hat{b}_{N}^{\dagger}|\emptyset\rangle$, where $|\emptyset\rangle$ denotes the vacuum state. For $\hat{H}_{2}$, the sublattice B in unit cell $n$ is coupled to the sublattice A in unit cell $n+2$ for n=1,,N2n=1,...,N-2. The sublattices A in unit cells $1,2$ and B in unit cells N1,NN-1,N are thus isolated from other sites, making them available to host four edge zero modes $\hat{a}_{1}^{\dagger}|\emptyset\rangle$, $\hat{a}_{2}^{\dagger}|\emptyset\rangle$, $\hat{b}_{N-1}^{\dagger}|\emptyset\rangle$ and $\hat{b}_{N}^{\dagger}|\emptyset\rangle$. It is also clear that the topological origin of zero-energy edge modes in a single $\alpha$ chain does not depend on whether the system is Hermitian or non-Hermitian. Note that for the $\alpha=1$ and $\alpha=2$ chains, their edge zero modes are robust to any perturbations (such as disorder) added to the hopping amplitudes $(J_{\alpha}^{\rm L},J_{\alpha}^{\rm R})$ over different unit cells, as these perturbations could not couple the isolated edge zero modes of these single $\alpha$ chains to bulk states. The topological degeneracy of these edge zero modes is protected the sublattice symmetry of the bulk.
Appendix C Calculation of the edge states
@@
In lattice representation, the Hamiltonian of NHSSH $(0,1)$ chain in Sec. III.2 takes the form
@@
Following Ref. [25], we apply a similarity transformation $\hat{S}$ to $\hat{H}_{01}$ under the OBC, where
@@
We choose $\varrho$ to ensure the left and right intracell hoppings to have equal magnitudes, which means that $|\rho|=\sqrt{|J+\gamma|/|J-\gamma|}$. Let us consider a chain with unit cell indices n=1,2,,Nn=1,2,...,N and take the OBC at both ends. A zero-energy eigenmode of $\hat{H}^{\prime}_{01}$ satisfies $\hat{H}^{\prime}_{01}|\psi_{0}\rangle=0$. Expanding $|\psi_{0}\rangle$ in the lattice representation as $|\psi_{0}\rangle=\sum_{m}(A_{m}\hat{a}_{m}^{\dagger}+B_{m}\hat{b}_{m}^{\dagger})|\emptyset\rangle$ and applying $\hat{H}^{\prime}_{01}$ to it, we get the iteration equations of wave amplitudes $(A_{n},B_{n})$, i.e.,
$\varrho^{-1}(J+\gamma)B_{1}=0,\qquad\varrho(J-\gamma)A_{N}=0,$ (37)
$\varrho(J-\gamma)A_{n}+A_{n+1}=0,\qquad n=1,...,N-1,$ (38)
$\varrho^{-1}(J+\gamma)B_{n+1}+B_{n}=0,\qquad n=1,...,N-1.$ (39)
@@
$|\psi_{0}^{{\rm R}}\rangle=$ $\sum_{n=1}^{N}[-(J+\gamma)/\varrho]^{N-n}\hat{b}_{n}^{\dagger}|\emptyset\rangle.$ (40)
It is clear that the states $|\psi_{0}^{{\rm L}}\rangle$ and $|\psi_{0}^{{\rm R}}\rangle$ represent left-localized and right-localized edge zero modes if and only if $|\varrho(J-\gamma)|<1$ and $|(J+\gamma)/\varrho|<1$, respectively. Both these conditions yield $|J^{2}-\gamma^{2}|<1$, which is exactly the condition for us to have topologically nontrivial gapped phases with the winding number $W=1$ and the number of zero-energy edge modes $N_{0}=2$, as reported in Eqs. (19) and (20). The bulk-edge correspondence of our NHSSH $(0,1)$ chain is thus proved. Importantly, we find no zero-energy edge modes when $J^{2}-\gamma^{2}=\pm 1$, i.e., at the critical points of the system under OBC. This observation confirms that the gapless phase boundary of the NHSSH $(0,1)$ chain is indeed topologically trivial from the perspective of critical edge states.
We next consider the NHSSH $(1,2)$ chain in Sec. III.2, whose Hamiltonian takes the form
@@
Under a similarity transformation
@@
$|\varphi_{0}^{{\rm L,1}}\rangle=\hat{a}_{1}^{\dagger}|\emptyset\rangle,\qquad|\varphi_{0}^{{\rm R,1}}\rangle=\hat{b}_{N}^{\dagger}|\emptyset\rangle,$ (44)
which form a pair of degenerate edge modes at the left and right ends of the lattice. To find other possible solutions, we again expand the zero mode in the lattice space as $|\varphi_{0}\rangle=\sum_{m}(C_{m}\hat{a}_{m}^{\dagger}+D_{m}\hat{b}_{m}^{\dagger})|\emptyset\rangle$ and apply the Hamiltonian $\hat{H}^{\prime}_{12}$ to it. The resulting zero-energy solutions in the limit $N\rightarrow\infty$ are given by
@@
$|\varphi_{0}^{{\rm R,2}}\rangle=$ $\sum_{n=1}^{N-1}[-(J+\gamma)/\varrho]^{N-n-1}\hat{b}_{n}^{\dagger}|\emptyset\rangle.$ (45)